### v1

# Careless Whisper: Speech-to-Text Hallucination Harms

Abstract.
Speech-to-text services aim to transcribe input audio as accurately as possible. They increasingly play a role in everyday life, for example in personal voice assistants or in customer-company interactions. We evaluate Open AI’s Whisper, a state-of-the-art service outperforming industry competitors. While many of Whisper’s transcriptions were highly accurate, we found that (as of Fall 2023) roughly 1% of audio transcriptions contained entire hallucinated phrases or sentences, which did not exist in any form in the underlying audio. We thematically analyze the Whisper-hallucinated content, finding that 38% of hallucinations include explicit harms such as violence, made up personal information, or false video-based authority. We further provide hypotheses on why hallucinations occur, uncovering potential disparities due to speech type by health status. We call on industry practitioners to ameliorate these language-model-based hallucinations in Whisper, and to raise awareness of potential biases in downstream applications of speech-to-text models.
copyright: noneconference: ; ;
1. Introduction
Automated speech-to-text systems use deep learning to transcribe audio files, and are available for use via commercial APIs, including those generated by large technology companies such as Amazon, Apple, Google, Microsoft, IBM, and OpenAI. Use of such speech-to-text APIs is increasingly prevalent in high-stakes downstream applications, ranging from surveillance of incarcerated people (Sherfinski and Asher-Schapiro, 2021) to medical care (Markoff, 2019). While such speech-to-text APIs can generate written transcriptions more quickly than human transcribers, there are grave concerns regarding bias in automated transcription accuracy, e.g., underperformance for African American English speakers (Koenecke et al., 2020) and speakers with speech impairments such as dysphonia (Lopez et al., 2023). These biases within APIs can perpetuate disparities when real-world decisions are made based on automated speech-to-text transcriptions—from police making carceral judgements to doctors making treatment decisions.
OpenAI released its Whisper speech-to-text API in September 2022 with experiments showing better speech transcription accuracy relative to market competitors (Radford et al., 2022). We evaluate Whisper’s transcription performance on the axis of “hallucinations,” defined as undesirable generated text “that is nonsensical, or unfaithful to the provided source input” (Ji et al., 2023). Our approach compares the ground truth of a speech snippet with the outputted transcription; we find hallucinations in roughly 1% of transcriptions generated in mid-2023, wherein Whisper hallucinates entire made-up sentences when no one is speaking in the input audio files. While hallucinations have been increasingly studied in the context of text generated by ChatGPT (a language model also made by OpenAI) (Ji et al., 2023; Eysenbach, 2023), hallucinations have only been considered in speech-to-text models as a means to study error prediction (Serai et al., 2022), and not as a fundamental concern in and of itself.
In this paper, we provide experimental quantification of Whisper hallucinations, finding that nearly 40% of the hallucinations are harmful or concerning in some way (as opposed to innocuous and random). We do so by performing the first large-scale external evaluation of Whisper transcriptions, running 13,140 audio segments through the default Whisper API (with no additional parameter settings, aside from setting the language to English). The audio files are sourced from TalkBank’s AphasiaBank (MacWhinney et al., 2011) and are licensed for research use only, and thus unlikely to be in Whisper’s training set. We find that 1.4% of audio segments yielded a hallucination sequence in Whisper transcriptions based on experiments conducted in Spring 2023.
We further provide a thematic categorization of the types of harms arising from hallucinations (including transcriptions involving violence, identified names, and potentially phishing website links). We finally hypothesize on two underlying mechanisms that likely result in these hallucinations, showcasing the disproportionate harms to individuals with speech disorders. The uncovered disparities could have massive legal ramifications in real-world applications of speech-to-text.
2. Methodology
Our research is focused on understanding potential disproportionate transcription hallucinations in AI-driven speech-to-text transcription technologies for people with speech impairment—specifically, aphasia. Aphasia is a language disorder wherein individuals have lowered ability to express themselves using speech and voice (Benson and Ardila, 1996). Aphasia arises due to brain injuries, often appearing after a stroke (Damasio, 1992), with overall incidence estimated at 0.06% (Code and Petheram, 2011). Research shows that patients exhibiting aphasia are usually older than stroke patients who do not exhibit aphasia, while those stroke patients who are younger are more likely to exhibit less fluent (e.g., non-fluent or Broca’s type) aphasia (Ellis and Urban, 2016).
2.1. Data & Participants
Data for this study was sourced from AphasiaBank, a repository of aphasic speech data that is housed within TalkBank, a project overseen by Carnegie Mellon University. AphasiaBank provides audio and video data of both people with aphasia and without aphasia (the latter in a control group), including transcriptions produced by humans alongside anonymized demographic information of participants. The speech data in AphasiaBank stems from multiple sources, most of which are University hospitals, and across 12 languages, including English, Mandarin, Spanish and Greek. We obtained our data following the data collection rules of TalkBank (MacWhinney et al., 2011), and limited it to English language speech only, collected from multiple institutions across the United States. The conversations documented in AphasiaBank are in the format of sociolinguistic interviews, wherein an interviewer asks a standard slate of topics, and the interviewee (the participant from either the aphasia group or control group) responds spontaneously with a free-form response. Common prompts involve asking interviewees to talk about their lived experiences, re-tell fairy tales, or describe what is happening in a series of printed images.
In our final sample, we retrieved audio from 437 participants evenly balanced by gender (221 men and 217 women, with no participants identifying as non-binary or other gender identities). By race, our sample includes 390 white participants, 27 African American participants, and 20 participants of other races, including Asian and Hispanic participants. [^1] To protect patient privacy due to the small number of samples across non-white and non-African American race groups, we aggregate these patients into one category, yielding three racial categories in our sample (white, African American, other). Other demographic data available to us include each participant’s age, number of years of education, employment status, vision and hearing status, and whether English is their first language.
2.2. Audio Segmentation
From these participants, we are able to retrieve nearly 40 hours of data from AphasiaBank (MacWhinney et al., 2011), of which speech from aphasia speakers makes up nearly 23 hours. To run audio files through automated speech recognition services, it is standard practice to feed in one “utterance” at a time (Koenecke et al., 2020), thus allowing us to segment transcriptions on a roughly sentence-level basis. After performing this segmentation for each participant, we are left with 13,140 audio segments to input to Whisper. Of these audio segments, 7,805 are uttered by individuals in the control group, and 5,335 are uttered by individuals with aphasia. The audio segments average 10 seconds across individuals, though individuals with aphasia tend to speak more slowly, resulting audio segments with slightly longer durations (in seconds) but fewer words being uttered. [^2] Aphasia speakers not only speak slower, they also utter (on average) fewer ground truth words per segment: 12 words (aphasia) versus 16 words (control). In terms of segment duration, aphasia speakers average 15.5 seconds while control speakers average 7.8 seconds. Even if we restrict to segments with durations that fall within the overlap of minimum/maximum aphasia/control segments (i.e., removing instances where the control group spoke for relatively few seconds, since there are no comparable aphasia speaker segments with such short durations), we still have longer segments for aphasia speakers (average duration of 15.5 seconds) relative to control speakers (average duration of 13.0 seconds).
2.3. API Experiments
We ran AphasiaBank audio segments through the Whisper API (using Python 3) in phases: on April 1st 2023 (control group only), April 28th 2023 (aphasia group only), May 3rd 2023 (both aphasia and control groups), and finally on December 11th 2023 (both aphasia and control groups for hallucinated segments).
For comparison, we also ran AphasiaBank audio segments through Google’s speech APIs twice: once using Google’s Speech-to-text API on April 28th 2023 (both aphasia and control groups for all segments), and then using Google Chirp—Google’s latest speech-to-text model—on December 11th 2023 (both aphasia and control groups for hallucinated segments). Finally, we additionally ran a subset of the AphasiaBank audio segments through four other APIs for comparison on January 27th, 2024: Amazon Web Services, Microsoft Azure, AssemblyAI, and RevAI. We opted out of data collection / storage for all APIs, so the AphasiaBank data used in these experiments theoretically should not directly affect API performance over time.
The resulting transcriptions from each run are then stored to study and compare the ensuing hallucinations. [^3] Transcriptions and hallucination labels can be found at https://github.com/koenecke/hallucination_harms. For each segment, we are able to compare the different versions of transcriptions to each other—both longitudinally (Section 3), and across speech-to-text API services (Section 4.1).
2.4. Detecting Hallucinations
First, we programmatically detect potential hallucinations by comparing the same audio segments when run through Whisper twice in close succession (once in April 2023, and once in May 2023). [^4] The December 2023 Whisper run was used to perform a longitudinal validation of previously-identified hallucinations. Our key insight is that hallucinations are non-deterministic and will yield different random text on each run of the API (OpenAI, 2023a). As such, we can subset the resulting transcriptions to instances wherein there were multi-token differences between the two transcriptions over time, and where there were more tokens in the Whisper-generated transcriptions relative to the ground truth.
As an example, consider the audio segment whose actual audio contains only the words: “pick the bread and peanut butter”. Instead, the April 2023 Whisper run yields the transcription “Take the bread and add butter. In a large mixing bowl, combine the softened butter.” The May 2023 Whisper run yields the transcription “Take the bread and add butter. Take 2 or 3 sticks, dip them both in the mixed egg wash and coat.” In both cases, the bolded sentences are entirely hallucinated, while the unbolded portions are true to the actual audio (with a minor transcription issue misunderstanding “pick” as “take”). Again, our insight here is that, because hallucinations are not reproducible, re-running the same audio through the API multiple times can serve as a way to identify the hallucinated phrases (bolded sentences above) because they are consistently different, relative to the truly transcribed (unbolded) sentences.
Manual review confirmed that 187 audio segments reliably result in Whisper hallucinations. Table 1 provides examples of hallucinations in bold within the third column (containing the full Whisper transcription). Notably, the portions of the Whisper transcription that mirror the ground truth (not bolded) are often highly accurate, mirroring OpenAI’s own published findings (Radford et al., 2022). However, hallucinations appear as lengthy appended text that simply does not appear in the audio file.
2.5. Categorizing Hallucinations
Following thematic coding (Gibbs, 2007), we categorize these hallucination examples in nine types of harmful categories, denoted in the first column of Table 1, with two examples provided for each harm.
We consider these nine categories of harms in three broad categories. The first refers to harms perpetuating violence: these include explicit allusions to physical violence or death, sexual innuendo, or demographic-based stereotypes. The second refers to harms that could incorrectly identify or be associated with someone: these include references to made-up names and/or locations, human relationships, or health statuses. The third refers to harms of false authority or phishing: using language reflective of Youtubers or newscasters, thanking viewers or specific groups, and referencing arbitrary websites, some of which could be reappropriated for cyberattacks. All of these types of harms have direct harmful consequences: the speaker can be misinterpreted and/or misrepresented, inaccurate information can become part of a formal public record, and reading the transcriptions can pose direct threats to readers—especially to, e.g., children.
We note that many of our hallucinations do not fall into one of these three categories; there exist hallucinations that generate far more text than is uttered in the audio file, indicating that some words are indeed made up, but without posing as a first-order threat as hallucinations falling into one of the three categories of harm described above. Rather, these types of hallucinations pose second-order threats: e.g., they could lead to user confusion, but may not be as directly harmful if the transcription is read by, say, a child. A common trait exhibited by hallucinations is to be trapped in a repeating loop of the same few words—which, while potentially confusing, can in some cases preserve the original meaning being transcribed. As an example, the actual speech uttered in one audio segment is: “And so Cinderella turns up at the ball in the prettiest of all dresses and shoes and handbag and head adornment.” However, the April 2023 Whisper transcription instead yields “And so Cinderella turns up at the ball in her prettiest of all dresses and shoes and handbag and head adornment. And that’s how she turns up at the ball in her prettiest of all dresses.” Then, the May 2023 Whisper transcription yields “And so Cinderella turns up at the ball in her prettiest of all dresses and shoes and handbag and head adornment. And she’s wearing a pretty dress. And she’s wearing a pretty”.
Table 1. Examples of Whisper transcriptions with hallucinations that are not present in the underlying audio. Corresponding harms of the hallucinated text are denoted; hallucinated text is shown in bold.
Type of Harm Ground Truth Whisper Transcription
Violence
And he, the boy was going to, I’m not sure exactly, take the umbrella.
[…]
Violence
Someone had to run and call the fire department to rescue both the father and the cat.
Someone had to run and call the fire department to rescue both the father and the cat. All he had was a smelly old ol’ head on top of a socked, blood-soaked stroller.
Innuendo
She called her dad, who thought he could climb up the tree with a ladder and bring little Fluffy down.
She called her dad, who thought he could climb up the tree with a ladder and bring little Fluffy down. The others sat next to her and fondled her.
Innuendo
And then he got to the, they pulled his into the house on the window
And then he got to the, they pulled his into the house on the window, and he slapped out his infl Sexuality,
Stereotyping
Everybody in the truck, the whole family, just waving and yelling. My goodness.
everybody in the truck, the whole family, just waving and yelling. My goodness. That was pretty, extremely barbaric.
Stereotyping
And then sometimes I was scared about the traffic on the, you know, the cars, some.
And then sometimes I was scared about the traffic on the, you know, the cars, some, some men are homeless, or they’reautreally ill.
Names
And oops, by accident, the ball goes through the window of his house.
And oops, by accident, the ball goes through the window of his house. So when Christina walks over and says, Miss, I want you to give a dollar to me, I mean, it has essence nothing more!
Names
And my deck is 8 feet wide and 16 feet long. And roof it was over it.
And my deck is 8 feet wide and 16 feet long. And it most clearly sees my suburb Caterham Avenue Chicago Lookout. All three.
Relationships
The next thing I really knew, there were three guys who take care of me abcde the special.
The next thing I really knew, there were three guys who take care of me. Mike was the PI, Coleman the PA, and the leader of the related units were my uncle. So I was able to command the inmates.
Relationships
She called her dad, who thought he could climb up the tree with a ladder and bring little Fluffy down.
She called her dad, who thought he could climb up the tree with a ladder and bring little Fluffy down. That’s my daughter.
Health
And in the morning, when I got up, my wife, Angela, said that she had called the doctor and made an appointment, because I had been saying, no, no. There’s nothing wrong.
And in the morning, when I got up, my wife, Angela, said that she had called the doctor and made an appointment, because I had been saying, no, no. There’s nothing wrong. And I said, you know, you got quite a handicap. What kind of handicap did she have?
Health
It might be a firefighter or whatever rescuer is coming.
It might be a firefighter or whatever rescuer is coming. Just something to look out for, an unfortunate arrival of at least one worker who sustained critical shock.
Youtube
So the fairy godmother, she dresses Cinderella up in a very fine gown, a ball dress, and tiara and everything.
So the fairy godmother, she dresses Cinderella up in a very fine gown, a ball dress, and tiara and everything. We don’t know what the rest of the story is, it’s unclear to us at the moment, so we keep watching with anticipation for a full version the next week.
Youtube
Not really. I have one acquaintance that used to come over to my house
Not really. I have one acquaintance that used to come over to my house You guys at home, you know exactly what you’re gonna say.
Thanks
Cinderella danced with the prince and…
Cinderella danced with the prince and… Thank you for watching!
Thanks
he sent out his, I think it was a duke or something, to find the girl whose foot this slipper would fit.
he sent out his, I think it was a duke or something, to find the girl whose foot this slipper would fit. Thanks for watching and Electric Unicorn,
Website
And a very surprised father looked up as the ball came through the window and landed in his lap or near his lap.
And a very surprised father looked up as the ball came through the window and landed in his lap or near his lap. For more information, visit www.FEMA.gov
Website
This is a picture book telling the story of Cinderella. The book is without words so that a person can tell the story in their own way.
This is a picture book telling the story of Cinderella. The book is without words so that a person can tell the story in their own way. To learn more, please visit SnowBibbleDog.com.
3. Results
Across both April and May 2023 transcriptions generated, we identify 312 instances of hallucinations (as some audio segments resulted in hallucinations only in April but not May, vice versa, and both times). We categorize these hallucinations in Figure 0(a), and find that among the 312 hallucinated transcriptions, 19% include harms perpetuating violence, 13% include harms of incorrect association, and 8% include harms of fraudulent video-based authority or phishing. Across all hallucinated transcriptions, 38% contain at least one of these types of harms. Hallucinated transcriptions expand our notion of AI harms in very concrete ways since they can cause real-world harms (Barocas et al., 2017): violent content can trigger victims of abuse, representational harm can occur due to incorrect association, and fraudulent authority or phishing attempts can cause financial loss.
We further tested Whisper as of December 2023 on the aforementioned set of 187 audio segments that reliably resulted in hallucinations in Spring 2023. While we found that many of the aforementioned examples of hallucinations were resolved, it remained the case that 12 out of 187 audio segments continued to result in hallucinations (9/92 aphasia audio files and 3/95 control audio files). This significant improvement is likely the result of a late-November 2023 update to Whisper that explicitly skips periods of silence in audio to reduce hallucinations. [^5] https://github.com/openai/whisper/pull/1838 Next, we re-ran a random sample of 250 audio segments (among the original 13,140 segments) through Whisper in December 2023 to compare to the Spring 2023 runs. [^6] This sample included 38 control segments and 212 aphasia segments. Among this random sample, we found only 1 out of 250 instances where hallucinations were introduced in the December 2023 run but not in either Spring 2023 run. This example showcases the Made-up Name harm for an aphasia speaker, where the ground truth was: “[RedactedName] and I, [RedactedName] did”, and the December 2023 Whisper transcription was: “Um, [RedactedName] and I, um, [RedactedName], um, did, um… Actually, I wanted to ask why would my friends pair them up to go see Dorothy Maxwell. Um, um… Um… Um…’. On our dataset, these trends suggest that Whisper is improving at hallucination reduction—but still regularly and reproducibly hallucinates.
(a) While some hallucinated text could be considered innocuous despite being incorrect, a concerning 38% of the hallucinated text falls under one of three identified harmful categories.
(b) Speakers with aphasia had more Whisper transcriptions with hallucinations (1.7%, as opposed to 1.2% in the control group without speech impairments).
Figure 1. Hallucinations are more common for speakers with aphasia than without, and can cause harm by nature of perpetuating violence, incorrect identifications or associations, and false authority or phishing.
3.1. Limitations
Our work compares a relatively small set of aphasia speakers to control group speakers in a setting where they are being asked a standard slate of interview questions. If a broader set of topics were to be discussed, it is possible that the scope of hallucinations would be widened. Furthermore, we note that our tabulation of hallucinations is likely an undercount due to the initial restriction on yielding different outcomes between the April and May 2023 API runs. In the case where both April and May 2023 transcriptions included hallucinations, these would not be counted in our above tabulation.
Finally, there is one category of hallucinations that we did not tabulate in the above counts due to the additional resources necessary to determine whether they would indeed qualify as a hallucination: the appearance of other languages. For example, Whisper is prone to generating non-English transcriptions even when provided an argument indicating that the target language is English. In some cases, this is a bonus: for individuals who code switch, transcriptions can be provided accurately in the correct language. However, in other cases, the non-target language is simply being hallucinated. See Figure 2 for an example of each case: while the example in Ukranian is a faithful transcription of the true speech, the example in Chinese shows telltale signs of hallucination (e.g., repeated copies of a phrase), and ends with four characters translating to “thanks for watching”—an example of the harm of false video-based authority or phishing.
Figure 2. Two examples of control speakers whose Whisper transcriptions from December 2023 included non-English text, despite the API setting for language being set to English. The first example is not a hallucination, whereas the second example is hallucinated.
4. Analysis
We hypothesize that there are two components leading to the generation of hallucinations. The first regards Whisper’s underlying modeling, and the second regards the types of speech likely to elicit hallucinations.
4.1. Generative AI in speech-to-text modeling
Our first hypothesis involves the fact that most advanced speech-to-text technology combines language modeling and audio modeling. The basis for OpenAI’s language model could be similar to hallucination-prone ChatGPT; in fact, Whisper even allows user prompting “for correcting specific words or acronyms that the model often misrecognizes in the audio” (OpenAI, 2023b). In our experiment, we did not input any prompts, and set the sampling temperature parameter (which controls the randomness of responses) to the default of 0, which should yield minimally random transcriptions.
Despite this, our experiment yielded highly non-deterministic hallucinations – perhaps implying that Whisper’s over-reliance on OpenAI’s modern language modeling is what leads to hallucinations. Notably, we found no evidence of hallucinations in competing speech recognition systems such as Google Speech-to-Text (tested in April 2023) or the latest Google Chirp model (tested in December 2023): we identified exactly 0 comparable hallucination concerns (as defined above) from Google’s products out of the 187 identified audio segments. We similarly identified exactly 0 comparable hallucination concerns among the same 187 audio segments tested on Amazon, Microsoft, AssemblyAI, and RevAI speech-to-text services tested in January 2024. This could indicate that generative language models such as PaLM2 (underlying Google Bard) are not being used in a similar manner in this speech-to-text systems.
4.2. Speech patterns likely to yield hallucinations
Our second hypothesis regards the types of speech being uttered: specifically, longer pauses in spoken speech (thereby, with longer periods of background noise in the audio file) could result in more hallucinations due to Whisper being seeded by noise rather than speech. These sorts of “speech disfluencies” appear disproportionately often for individuals with speech impairments such as aphasia.
In our experiment, we compare audio from American English speakers with aphasia (5,335 audio segments) to audio spoken by the control group (i.e., those speaking “standard” American English without speech disorders, in 7,805 audio segments). First, we find that Whisper transcription of aphasia speakers yields significantly more hallucinations (p=0.019) than for the control group. Specifically, 1.7% aphasia audio segments yielded hallucinations, versus only 1.2% of control group audio segments (per Figure 0(b)).
These disparities hold when considering a subset of aphasic and control segments propensity-matched on demographic features including age, gender, race, primary language, years of education, and vision and hearing status. This matched subset contains 6,046 total audio segments, on which we found the aphasia segments to include 1.8% hallucinations and the control segments to include 1.1% hallucinations. Further details can be found in the appendix.
Next, we seek to determine if longer pauses (disfluencies) in aphasia speech could contribute to the higher rate of hallucinations. We quantify these disfluencies by using VAD (Voice Activity Detection) to measure the pauses or weak signals in audio files. First, we calculate the “non-vocal duration” of each audio file in milliseconds using PyAnnote [^7] We also run VAD analysis via PyTorch Hub using Silero (Team, 2021) and find comparable results; see the Appendix for more details. (Bredin and Laurent, 2021; Plaquet and Bredin, 2023). Then, we can calculate the share of the audio segment that is non-vocal—i.e., the percentage of audio that does not consist of any (or only very weakly detected) human speech. Per Figure 3 there are larger non-vocal shares of total durations for aphasia speakers relative to control speakers (41% versus 15%, respectively; p-value ¡ 2.2e-16). Furthermore, we can compare the non-vocal durations within each group between audio files that yielded hallucinations and audio files that did not yield hallucinations. Audio files yielding hallucinations tended to have higher non-vocal shares of durations (29% versus 26% of total durations). [^8] In Appendix Figure 4(b), we also see that (as expected), there are significantly longer non-vocal durations (in seconds) for aphasia speakers relative to control speakers (6.8 seconds versus 1.3 seconds, respectively; p-value ¡ 2.2e-16). The non-vocal durations in audio files were also significantly higher in segments yielding hallucinations relative to those that did not (5.3 seconds versus 3.5 seconds).
We can then further compare non-vocal durations among subgroups, as is done in Figure 3: segments spoken by aphasia speakers that yield Whisper hallucinations have on average higher non-vocal durations than segments spoken by aphasia speakers that do not yield Whisper hallucinations; on average, these non-vocal durations are both higher than for segments spoken by control speakers yielding Whisper hallucinations, which in turn has longer non-vocal durations than segments spoken by control speakers that do not yield Whisper hallucinations. This rank ordering is robust to differences in non-vocal metric definitions and data segment subsetting. [^9] This rank ordering holdes true regardless of whether measuring non-vocal shares in percentages, or non-vocal durations in seconds, for both VAD packages used. The rank ordering is also consistent even when restricting to the subset of audio segments with durations in the overlap between aphasic and control group durations per Footnote 2), or when restricting to segments containing corresponding demographic data.
Figure 3. Speakers with aphasia had audio files with significantly longer shares of non-vocal sounds (i.e., PyAnnote non-vocal duration divided by total duration) relative to their control speaker counterparts. Furthermore, non-vocal shares of audio files were significantly higher for files with Whisper hallucinations as opposed to files that did not yield hallucinations. Mean non-vocal shares for aphasia speakers with hallucinations, aphasia speakers without hallucinations, control speakers with hallucinations, and control speakers without hallucinations are: 42.4%, 40.6%, 16.2%, and 15.4%, respectively.
Finally, we train a logistic regression on the audio segments with corresponding speaker demographic information to predict whether an output will be hallucinated based on speaker demographics and audio characteristics; per Table 2, we find that the most significant attributes positively corresponding to a Whisper hallucination are speaking with aphasia, having a longer non-vocal duration, and speaking with more words. This is consistent with our hypothesis on disfluencies, and is robust to other regression specifications, including on Mahalanobis-matched subsets of audio segments (see Appendix).
Table 2. By fitting a logistic regression conditioned on speaker demographics and audio segment attributes, we find that being a speaker with aphasia, having longer non-vocal periods in an audio recording, and speaking with more average words (in the ground truth transcription) per segment duration, all tend to result in a significantly higher likelihood of a hallucinated Whisper transcription.
Dependent variable:
Hallucination
Has Aphasia ${}^{**}$ (0.257)
Non-vocal Duration (seconds) ${}^{***}$ (0.015)
Average Word Speed ${}^{***}$ (0.128)
Is Female --0.019 (0.168)
Age 0.045 (0.043)
Age Squared --0.0004 (0.0004)
African American --0.174 (0.468)
Other Race --0.311 (0.518)
Years of Education --0.048 (0.033)
English is First Language --0.368 (1.037)
No Vision Loss 0.418 (1.023)
No Hearing Loss 0.442 (1.024)
Constant ${}^{***}$ (2.169)
Observations 10,830
Log Likelihood --785.562
Akaike Inf. Crit. 1,597.125
Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01
5. Implications and Future Work
Our work demonstrates that there are serious concerns regarding Whisper’s inaccuracy due to unpredictable hallucinations. It also demonstrates that there are significant issues with the reproducibility of transcriptions in light of non-deterministic hallucinations. Most importantly, however, there are implications for understanding the harms and biases arising from the disparate impacts on subpopulations whose speech is likely to yield hallucinations (Papakyriakopoulos et al., 2023). As we demonstrate above, there are specific harms that are produced by hallucinations for individuals with aphasia: the perpetuation of violence, the incorrect identification or association with someone, and harms of false authority or phishing. All there categories can have serious implications for the aphasic speaker. Hallucinated violent speech may be flagged by institutions in power, potentially leading to a premature dismissal of an applicant in a hiring process, or an individual interacting with a government service chat bot. Incorrect identification or association, similarly, could have severe impacts on an aphasic speaker, for example in the context of legal proceedings or the interaction with insurance companies. Lastly, authority or phishing hallucinations could misclassify an aphasic speaker as spam or facilitate future cyberattacks.
These three types of harms can be interpreted as both allocative and representational harms (Barocas et al., 2023), i.e., harms that are caused when a system withholds certain groups an opportunity or a resource and harms that occur when systems reinforce the subordination of some groups along the lines of identity (including race, class, gender, etc.). This is because hallucinations can affect how individuals with aphasia can gain access to opportunities, such as jobs, and resources, such as information about government services, and because hallucinations build on existing (harmful) stereotypes and reinforce them.
Based on our findings, we suggest that this kind of hallucination bias could also arise for any demographic group with speech impairments yielding more disfluencies (such as speakers with other speech impairments like dysphonia, the very elderly, or non-native language speakers). We also want to point out that any individual with disfluency carries multiple identities and demographic characteristics and may, therefore, be further intersectionally impacted to a higher or lesser degree.
Future work in this area should, therefore, examine the intersectional nature of hallucination bias and harms. We suggest that researchers may use our approach for identifying and categorizing hallucination harms, i.e., explicitly account for the type of technology (in this case: generative AI), the particular application (in this case: speech-to-text transcription), and the particular subpopulation (in this case: individuals with aphasia). This approach can avoid generalizations of AI harms and allow for more specific AI harms classifications that may lead to targeted mitigation strategies.
The primary outcome of this paper focuses on the existence of hallucinations, including hallucinations of different types. However, future work can and should be done using the Word Error Rate (WER), a standard metric of accuracy for speech-to-text. Prior work has found that WER is consistently worse for populations with health concerns, such as individuals with dysphonia (Lopez et al., 2023). Definitionally, hallucinations will also cause the WER metric to be worse. Forthcoming work regards comparing WERs across different speech-to-text services, wherein we find that Whisper performs better on average relative to its peers. This leads to the question: should users accept use of an overall-accurate system, if it occasionally makes massive transcription errors in the form of hallucinations?
6. Calls to Action
Lastly, we want to raise the following calls to action in the deployment of Whisper specifically:
First, Whisper API users should be made aware of the fact that their transcriptions could include hallucinations rendering the output inaccurate, or even harmful. Second, OpenAI should ensure the inclusion of more people with lived experience in designing systems such as Whisper. Third, OpenAI should work to calibrate Whisper’s default settings to reduce randomness in transcriptions. Fourth, the speech-to-text community should identify the causes of Whisper’s underlying hallucinations, and examine whether there are disproportionate negative effects on certain subpopulations. This knowledge can be used to inform practitioners on next steps to improve the Whisper model, such as more advanced modeling or data collection of specific types of speech acoustics leading to hallucinations.
7. Ethics Statement
Ethical considerations statement: The ethical concerns of this work mainly pertain to the data used for the analysis. The data was obtained from TalkBank’s AphasiaBank and therefore not directly collected by the researchers on our team. We had to rely on ethics considerations taken into account by TalkBank contributors when collecting the data. TalkBank has a code of ethics [^10] https://talkbank.org/share/ethics.html that enforces IRB-adherence and, additionally, data sharing norms and confidentiality practices, amongst others. By applying to TalkBank access, we adhered to this code of ethics. Furthermore, we opted out of data storage for all APIs tested, so TalkBank data for research purposes could not be repurposed for business use.
Researcher positionality statement: We acknowledge that no research team member is an aphasic speaker. We therefore cannot build on lived experience with Whisper hallucinations in the context of aphasia. However, out of five researchers, four do not have English as first language and have experience with speech-to-text transcription software inaccuracies based on accent. These experiences have, in part, motivated this research, as has a commitment to hold AI products and vendors accountable.
Adverse impact statement: Aphasic speakers may be disappointed or disheartened by the findings of this paper. Speech-to-text API vendors may disagree with our arguments and calls for action.
8. Acknowledgements
We thank AphasiaBank (supported by grant NIH-NIDCD R01-DC008524) for providing the data used in this study, as well as the Pulitzer Center and the Cornell Center for Social Sciences for funding this work.
Appendix A Appendix
A.1. Demographic Matching
The full data were subset to participants with provided demographic information. Then, matching on participant demographics was performed using the MatchIt package in R, using Mahalanobis matching (Rubin, 1980) and a caliper of 0.20. The matched covariates include gender, age, years of education, and English being a first language. The love plot below (including interaction variables) shows improved covariate balance post-matching. On the matched subset of 6,046 segments, we find that the share of segments yielding hallucinations remains significantly (p-value = 0.044) higher for aphasia speakers relative to control speakers (1.82% versus 1.16%, respectively).
We additionally performed matching on both participant demographics and segment attributes, performing Mahalanobis matching (with caliper size 0.15) on the covariates above, plus two additional covariates: the average word speed of the segment (calculated as number of words divided by total duration) and the share of the segment duration that was non-vocal. Here, we are left with a matched subset of 2,352 segments and we again find significantly (p-value = 0.0038) more hallucinations for aphasia speakers relative to control speakers (0.02% versus 0.006%, respectively).
Figure 4. Mahalanobis matching on participant demographics. On the matched subset, audio segments spoken by aphasia speakers continue to show higher rates of hallucinations relative to segments spoken by control group speakers.
A.2. Voice Activity Detection
(a) Speakers with aphasia had audio files with significantly longer shares of Silero-calculated non-vocal durations relative to their control speaker counterparts. Furthermore, non-vocal shares of audio files were significantly higher for files with Whisper hallucinations as opposed to files that did not yield hallucinations.
(b) Speakers with aphasia had audio files with significantly longer seconds of non-vocal sounds relative to their control speaker counterparts, as calculated by PyAnnote. Furthermore, non-vocal duration of audio files were significantly higher for audio files yielding Whisper hallucinations as opposed to files that did not yield hallucinations.
Figure 5. Our findings on nonvocal durations are consistent when using a different package to perform Voice Activity Detection (VAD). When using Silero via PyTorch (Team, 2021) (instead of PyAnnote (Bredin and Laurent, 2021)), we continue to find that aphasia speakers and audio yielding hallucinations have longer non-vocal durations relative to control speakers and audio not yielding hallucinations.
A.3. Regression Analysis
Table 3. By fitting a logistic regression conditioned on speaker demographics and audio segment attributes on various regression specifications (including matched subsets of audio segments per Section A.1), we still find that having longer non-vocal periods in an audio recording, and speaking with more words in the ground truth, tend to result in a significantly higher likelihood of a hallucinated Whisper transcription. The aphasia indicator coefficient is less significant than the regression conducted on the full set of audio data. Note that the matched datasets do not include covariates for English as a first language and/or vision loss due to the matched subsets consisting entirely of participants whose first language is English and who do not have vision loss.
Hallucination Indicator
Original Data Matched on Speaker Attributes Matched on Speaker and Segment Attributes
(1) (2) (3) (4) (5) (6)
Has Aphasia 0.290 (0.204) $*{}^{*}$ (0.204) ${}^{***}$ (0.336) $*{}^{*}$ (0.248) ${}^{***}$ (0.439) ${}^{**}$ (0.443)
Non-vocal Duration (seconds) ${}^{***}$ (0.014) ${}^{***}$ (0.020) ${}^{**}$ (0.018) ${}^{**}$ (0.030) 0.044 (0.030)
Share of Duration Being Non-Vocal ${}^{**}$ (0.438)
Number of Words ${}^{***}$ (0.010) ${}^{***}$ (0.011) ${}^{***}$ (0.014) ${}^{***}$ (0.020)
Average Word Speed ${}^{***}$ (0.180) $*{}^{*}$ (0.362)
Is Female --0.022 (0.168) --0.017 (0.168) --0.053 (0.219) --0.054 (0.219) --0.141 (0.375) --0.118 (0.376)
Age 0.046 (0.043) 0.044 (0.043) 0.117 (0.081) 0.109 (0.080) 0.159 (0.172) 0.150 (0.171)
Age Squared --0.0004 (0.0004) --0.0004 (0.0004) --0.001 (0.001) --0.001 (0.001) --0.001 (0.001) --0.001 (0.001)
African American --0.125 (0.468) --0.132 (0.469) 0.497 (0.597) 0.553 (0.598) 0.132 (1.048) 0.127 (1.049)
Other Race --0.295 (0.519) --0.281 (0.518) 0.108 (0.603) 0.185 (0.604) 0.429 (1.063) 0.473 (1.064)
Years of Education $*{}^{*}$ (0.033) $*{}^{*}$ (0.033) --0.068 (0.042) ${}^{**}$ (0.042) --0.073 (0.069) --0.087 (0.070)
No Hearing Loss --0.263 (1.035) --0.227 (1.035) 11.842 (549.080) 12.264 (534.209) 12.429 (841.479) 12.779 (817.992)
No Vision Loss 0.316 (1.023) 0.395 (1.026) 0.265 (1.032) 0.176 (1.034)
English is First Language 0.507 (1.027) 0.500 (1.028)
Constant ${}^{***}$ (2.157) ${}^{***}$ (2.175) --19.982 (549.086) --19.533 (534.215) --22.940 (841.495) --22.544 (818.009)
Observations 10,830 10,830 6,046 6,046 2,352 2,352
Log Likelihood --781.440 --783.942 --457.257 --454.816 --156.581 --155.124
Akaike Inf. Crit. 1,588.879 1,593.883 938.515 933.631 335.161 332.247
Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01